get remaining scl1 DL120G7's running



7 years ago
6 years ago


(Reporter: dustin, Assigned: mlarrain)



According to the spreadsheet, we have three HP's that weren't powered, and are now, 4873, 4874, and 4861

one that's broken, 4867

one that had a bogus toe tag, 2M2124012Y

and two that are having problems logging in, 4846 and 4876.

Let's get those either fixed up or sent back to HP for repairs, so that by the time we're ready to deploy these systems, we have 100% capacity.
colo-trip: --- → scl1
Assignee: server-ops-releng → mlarrain
I brought 4873, 4874, and 4861 online.  Check the spreadsheet for details as to what's left.

Comment 2

7 years ago
I have removed the power cables from the machines that aren't powering up. These need to be pulled and brought to MTV1 to go out for service.
Assignee: mlarrain → jwatkins
(In reply to Matthew Larrain[:digipengi] from comment #2)
> I have removed the power cables from the machines that aren't powering up.
> These need to be pulled and brought to MTV1 to go out for service.

Double check with Rich, but these should have the same warranty as all our other HPs, onsite repair included.
Matt, can you chase down service on these units and get someone onsite?  I'd like to have these fixed *before* we put these hosts into production.
Assignee: jwatkins → mlarrain

Comment 5

7 years ago
I had asked Jake to get me serials and such for them and I can have Rich send someone onsite.

Comment 7

7 years ago
2M2122007L pingable but hangs
2M2121035A Not pingable
Asset tag 4867 BROKEN
2M212400W9 Cannot connect
2M2122007F iLo password is incorrect/failed ssl login
2M212400W1 Cannot connect
2M2124012Y We have asset tag but no server's serial matches this tag

Need the serial for 4867 and need to find where the server that matches serial 

Comment 8

7 years ago
Adding Jake to grab the serial for 4867 and see if he can located the other random server.

Comment 9

7 years ago
Figured out where 2M2124012Y was and the serial for 4867 updated DHCP and having them rebooted.

Comment 10

7 years ago
Also 2M2121035A was fixed
I'm still unable to configure ilo on either of those (spreadsheet updated).

Comment 12

7 years ago
Contacted Rich and gave him info regarding what is wrong and the serial #'s. Hopefully we will have a tech out soon.

Comment 13

7 years ago
Called HP they want us to try resetting the system maint switch on the system board switch #2 or clear the nvram. 

Will try that and call back if needed, they said if its this then we will be looking at having the system board replaced. 

Case # is: 4669342709

Comment 14

7 years ago
I will be on site tomorrow to see about fixing these.

Comment 15

7 years ago
2 of these need power cords and 2 of them still need there nvram switch flipped. Will finish up on Monday/Tuesday.

Comment 16

7 years ago
These machines are all good to go minus two of them needing power cables still.
Last Resolved: 7 years ago
Resolution: --- → FIXED
That doesn't sound R/F to me :)
Resolution: FIXED → ---

Comment 18

7 years ago
Everything is in working order. they just need power cables.
The last 2 have been given power cables and powered on.
Just to be clear, I plugged in 4846 and 4847
dl120g7-r4-4846	2M212400W9	4846
dl120g7-r4-4847	2M212400W3	4847

Before I left, I also found this one to be without a pwr cable so I plugged it in too
dl120g7-r4-4839	2M2122007L	4839

Comment 21

7 years ago is up and running properly but 4846 and 4839 are still having issues. Running the XML file across them.

Comment 22

7 years ago
All these machines are up and running now :)
Last Resolved: 7 years ago7 years ago
Resolution: --- → FIXED
Component: Server Operations: RelEng → RelOps
Product: → Infrastructure & Operations
You need to log in before you can comment on or make changes to this bug.